# Boundary four-point connectivities of conformal loop ensembles

### 30 Mar 2026 v1

@@ 1 Introduction
The final step constitutes the main technical novelty of the paper; see Theorem 1.4. The third-order differential equation admits a three-dimensional space of solutions, corresponding to three distinct Frobenius series. For the link patterns $(12)(34)$ (resp. $(14)(23)$), the leading-order asymptotics as the cross-ratio tends to $1$ (resp. 0) are governed by boundary three-arm exponents, and thus allow for direct identifications with the higher-order Frobenius series. For the remaining link pattern $(1234)$, however, the leading order corresponds to the lowest-order Frobenius series, and it is not a priori clear whether and how the other two Frobenius series may appear as subleading terms. To resolve this ambiguity, we carry out a refined analysis of the subleading asymptotics, based on the observation that the subleading term of the CLE partition function defined in [MW18] with two wired boundary arcs decays rapidly. This enables us to determine the higher-order contributions and to uniquely identify each boundary connectivity with a specific solution to the third-order equation.
@@ 1.1 Boundary four-point connectivities of Bernoulli percolation
as $\lambda\to 0$. Here $A$ is the same constant as in Theorem 1.1.
@@ 1.2 Boundary four-point connectivities of CLE
Since (1.5) only has two singular points 0 and $1$, we know that the three Frobenius series $V_{0},V_{h},V_{3h+1}$ converge for all $\lambda\in(0,1)$. When $h$ and $3h+1$ are both non-integers, the expressions $1+O(\lambda)$ above are indeed power series of $\lambda$. When $h$ or $3h+1$ is integer (for $\kappa\in(4,8)$,this only occurs when $\kappa=6,\frac{16}{3},\frac{24}{5}$), there might be logarithmic terms in $O(\lambda)$, and we find explicit solutions in these cases separately. We refer readers to [Inc44, Chapter XVI] for more background on the Frobenius method.
@@
In [FLPW24, Conjecture 1.2], the authors formulate a precise conjecture for the limiting connection probabilities of general critical loop $O(n)$ models, which also give the conjectural connection probabilities for multichordal $\operatorname{CLE}_{\kappa}$ with $2N$ boundary arcs (see [AMY25]). In the regime $\kappa\in(4,8)$, this conjecture was proved for $n=\sqrt{2}$ (FK-Ising) [FPW24] and for $2N=4$ [MW18]. In principle, by taking $2N=8$ and shrinking four of the eight boundary arcs, one could obtain the CLE boundary Green’s functions (1.3) via renormalized limits of the corresponding conjectural connection probabilities. However, this does not seem practically feasible, since the explicit formulas for these conjectural connection probabilities involve the pure partition functions, whose constructions are highly nontrivial [FK15, KP16, Wu20, Zha24, FLPW24]. It is therefore unclear whether our results can be recovered in this way from these conjectural formulas.
@@ 1.3 Discussions
We now briefly discuss our proof strategy. The first step is to identify a precise correspondence between $\operatorname{CLE}_{\kappa}$ and the $\operatorname{SLE}_{\kappa}$ bubble measure for $\kappa\in(4,8)$. To achieve this, we show that the loop sampled from the counting measure on boundary-touching loops in a $\operatorname{CLE}_{\kappa}$ configuration is equal in law to the unrooted $\operatorname{SLE}_{\kappa}$ bubble measure, see Proposition 2.2. This identification allows us to express the boundary Green’s functions of $\operatorname{CLE}_{\kappa}$ (1.3) in terms of those of the $\operatorname{SLE}_{\kappa}$ bubble measure. Furthermore, $\operatorname{SLE}_{\kappa}$ bubble measure is also the weak limit of the chordal $\operatorname{SLE}_{\kappa}$ when its starting point approaches its end point. According to [FZ23], the boundary Green’s functions of chordal $\operatorname{SLE}_{\kappa}$ are finite and satisfy a martingale property as the chordal $\operatorname{SLE}_{\kappa}$ grows. Combined with their smoothness, the martingale property then yields a pair of second-order PDEs, known as BPZ equations. Establishing such smoothness is nontrivial and relies on Hörmander’s hypoellipticity; see Lemma 2.8, as inspired by [Dub15b, FZ23]. Then, we apply a fusion procedure to these PDEs to derive a third-order differential equation satisfied by boundary Green’s functions of the $\operatorname{SLE}_{\kappa}$ bubble, which corresponds to collapsing two marked boundary points. As mentioned before, this step follows from the framework in [Dub15a], which also appeared in recent works [KLPR25, LPR25] for specific values of $\kappa$. This proves Theorem 1.3.
@@
Acknowledgment. This work was supported by the National Natural Science Foundation of China (Grant No. 12526204) and National Key R&D Program of China (No. 2021YFA1002700).
@@ 2 SLE bubble measures and CLE boundary Green’s functions
We first review the background on the SLE bubble measure in Section 2.1, and then show the relation between boundary-touching CLE loops and the (unrooted) SLE bubble measure in Section 2.2. In Section 2.3, we relate the CLE boundary four-point connectivities (1.3) to the boundary Green’s functions of SLE bubble and of chordal SLE. In Section 2.4, we show the smoothness of these boundary Green’s functions, thus obtaining a pair of second-order PDEs satisfied by the boundary Green’s functions of chordal SLE, which will be the basis of the fusion procedure in Section 3.
@@ 2.2 Boundary-touching CLE loops and unrooted SLE bubble measure
Now, since $\rho$ is obtained from the counting measure on boundary-touching loops and every $z\in\mathbbm{H}$ is surrounded by at most one CLE loop, we have ${\bf 1}_{\ell\text{ surrounds }z}\rho(d\ell)dz={\bf 1}_{\ell\cap\mathbbm{R}\neq\emptyset}\Theta_{z}(d\ell)dz$, where $dz$ is the Lebesgue measure on $\mathbbm{H}$. Combined with the above result, we find ${\bf 1}_{\gamma\text{ surrounds }z}\mu^{\mathrm{Bub}}(d\gamma)dz=C{\bf 1}_{\ell\text{ surrounds }z}\rho(d\ell)dz$. The result then follows by deweighting the Lebesgue areas of the regions surrounded by the bubble (resp. loop) on both sides. ∎
@@ 2.3 CLE boundary Green’s functions
The aim of this section is to relate the $\operatorname{CLE}_{\kappa}$ boundary Green’s functions defined in (1.3) to the limit of boundary Green’s function of chordal $\operatorname{SLE}_{\kappa}$; see Proposition 2.7. Due to symmetry, it suffices to focus on $G^{(1234)}(x_{1},x_{2},x_{3},x_{4})$ and $G^{(12)(34)}(x_{1},x_{2},x_{3},x_{4})$. We refer readers to Appendix A for further background of $\operatorname{SLE}_{\kappa}$ boundary Green’s functions, including the various definitions appeared in the literature [Zha22, FZ23] and their equivalence to the definition in this paper.
@@
Let $\mathfrak{C}$} be the same constant as in Proposition 2.2. Then for any $n\geq 2$, we have
@@
where the integration is taken over $\mu_{\mathbbm{H},y_{1},y_{2}}^{\#}(d\eta)$. (2.12) is equivalent to the boundary Green’s function considered in [FZ23]; see Proposition A.2.
@@ 2.4 Smoothness and second-order PDEs
Note that the Lie brackets ${\bf X},{\bf Y}]=-\sum_{i=1}^{3}\frac{1}{x_{i}^{2}}\partial_{x_{i}}$ and ${\bf X},[{\bf X},{\bf Y}]]=2\sum_{i=1}^{3}\frac{1}{x_{i}^{3}}\partial_{x_{i}}$. Since $x_{i}$’s are mutually different on $U$, the Lie algebra generated by $\bf X$} and $\bf Y$} has rank $3$ on $U$. Thus, $\mathcal{L}$ satisfies the Hörmander’s condition222i.e. for two vector fields $\bf X$} and $\bf Y$}, ${\bf X},{\bf Y}$ and their iterated Lie brackets ${\bf X},{\bf Y}],[{\bf X},[{\bf X},{\bf Y}]]$ etc. span the whole tangent space at every point.. By [Bon69, Theorem 5.2], we can then define a Poisson operator $\mathcal{P}:C(\partial U)\to C^{\infty}(U)\cap C(\overline{U})$ for $\mathcal{L}$ such that for any $\omega\in C(\partial U)$, $(\mathcal{P}\omega)|_{\partial U}=\omega$ and $\mathcal{L}(\mathcal{P}\omega)=0$.
@@ 4 Identification of solutions
Here to the right side we use Lemma 4.4. Combined with $r_{3}^{-h}r_{4}^{-h}|\mathsf{m}_{r_{3},r_{4}}|\to G_{\mathbbm{H},x_{1},x_{2}}(x_{3},x_{4})$, the result then follows. ∎
@@
where $\tau$ and $\phi$ is defined in Corollary 4.6, and
@@
Here and after, $C>0$ is some constant depending on $x_{1},x_{3},x_{4}$ and can vary from line to line. To see (4.14), first note that for any compact hull $A$, $A\cap\mathbbm{R}=[a,b]$, $x>b$ such that ${\rm dist}(x,A)>d_{1}>0$, for the conformal map $\psi:\mathbbm{H}\setminus A\to\mathbbm{H}$ with $\psi(a)=0$, $\psi(b)=\infty$ and $\psi(x)=1$, we have $|\psi^{\prime}(x)|\leq 4d_{1}^{-1}$ by Koebe’s 1/4 theorem. By conformal covariance, this implies the boundary one-point Green’s function $G_{\mathbbm{H}\setminus A,a,b}(x)\leq Cd_{1}^{-h}$. Now suppose $b<y<x$, ${\rm dist}(y,A)>d_{2}>0$ and $|x-y|>d_{1}+d_{2}$. According to the martingale property of the boundary two-point Green’s function (see [FZ23, Theorem 4.1]), we have
@@
here $\mathbbm{E}_{x}$ is with respect to the conditional probability measure $\mathbbm{P}_{x}$ of the chordal $\operatorname{SLE}_{\kappa}$ curve $\eta$ on $\mathbbm{H}\setminus A$ from $a$ to $b$ conditioned to hit $x$ (we denote this hitting time by $\sigma$), and $U^{*}$ is the connected component of $\mathbbm{H}\setminus(A\cup\eta[0,\sigma])$ such that $b$ is on its boundary. Furthermore, following the estimate in [FZ23, Lemma 3.5], for $s\in(0,1)$, the probability that a chordal $\operatorname{SLE}_{\kappa}$ on $\mathbbm{H}$ from 0 to $\infty$ hits $\partial B(1,\frac{1}{10})$ after hitting $B(1,\frac{1}{10}s)$ and before disconnecting $1$ from $\infty$ is bounded above by $Cs^{2h}$. The domain Markov property of $\operatorname{SLE}_{\kappa}$ (see e.g. [FZ23, Eq.(3.27)]) then implies $\mathbbm{P}_{x}[{\rm dist}(U^{*},y)<rd]\leq Cr^{2h}$ for $r\in(0,\frac{1}{10})$. Consequently, we obtain $G_{\mathbbm{H}\setminus A,a,b}(x,y)\leq Cd_{1}^{-h}d_{2}^{-h}$, thus proving (4.14).
@@
Then (4.17) has three linearly independent solutions $V_{0}(\lambda)$, $V_{2}(\lambda)$ and $V_{\frac{1}{3}}(\lambda)$, which are defined in Section 1.1. See also [GV18, Appendix C]. In particular, $V_{0}(\lambda)=V_{0}(1-\lambda)$, and as $\lambda\to 0$,
@@
    They correspond to $G^{(14)(23)}$ and $G^{(12)(34)}$, respectively, and have the following probabilistic interpretation. Let $e_{1}$ (resp. $e_{2}$) be the union of two independent Brownian excursions from 0 to $\infty$ (resp. from $\lambda$ to $1$), and let $P(\lambda)$ be the probability that $e_{1}$ does not intersect with $e_{2}$. Then we expect $P(\lambda)\propto\frac{U_{1}(\lambda)}{\lambda^{4}}$. Namely, we have the conjectural formula
@@ 5 One-Bulk and two-boundary connectivities
By symmetry, the above PDE also holds when changing the place of $y_{1}$ and $y_{2}$. Here for convenience, we view $z$ and $\bar{z}$ as independent variables. Note that such equations are similar to those in Proposition 2.9, with the three marked points $(x_{1},h),(x_{2},h),(x_{3},h)$ replaced by $(x,h),(z,\frac{\alpha}{2}),(\bar{z},\frac{\alpha}{2})$. Consequently, we can repeat the fusion procedure in Section 3 in parallel. Namely, (5.2) implies a third order differential equation for $G(u,x,z)$. The resulting ODE [^5] Since $\lambda\in\mathbbm{C}$ takes value on the circle $\{\lambda\in\mathbbm{C}:|\lambda-1|=1\}$, we can view $\lambda$ as if it were a real variable. for $\Delta(\lambda)$ is

### 06 Apr 2026 v2

@@ 1 Introduction
The final step constitutes the main technical novelty of the paper; see Theorem 1.4. The third-order differential equation admits a three-dimensional space of solutions, corresponding to three distinct Frobenius series. For the link patterns $(12)(34)$ (resp. $(14)(23)$), the leading-order asymptotics as the cross-ratio tends to 11 (resp. 0) are governed by boundary three-arm exponents, and thus allow for direct identifications with the higher-order Frobenius series. For the remaining link pattern $(1234)$, however, the leading order corresponds to the lowest-order Frobenius series, and it is not a priori clear whether and how the other two Frobenius series may appear as subleading terms. To resolve this ambiguity, we carry out a refined analysis of the subleading asymptotics, based on the observation that the subleading term of the CLE partition function defined in [MW18] with two wired boundary arcs decays rapidly. This enables us to determine the higher-order contributions and to uniquely identify each boundary connectivity with a specific solution to the third-order equation.
@@ 1.1 Boundary four-point connectivities of Bernoulli percolation
as $\lambda\to 0$. Here AA is the same constant as in Theorem $1.$
@@ 1.2 Boundary four-point connectivities of CLE
Since (1.5) only has two singular points 0 and 11, we know that the three Frobenius series $V_{0},V_{h},V_{3h+1}$ converge for all $\lambda\in(0,1)$. When $h$ and $3h+1$ are both non-integers, the expressions $1+O(\lambda)$ above are indeed power series of $\lambda$. When $h$ or $3h+1$ is integer (for $\kappa\in(4,8)$, this only occurs when $\kappa=6,\frac{16}{3},\frac{24}{5}$), there might be logarithmic terms in $O(\lambda)$, and we find explicit solutions in these cases separately. We refer readers to [Inc44, Chapter XVI] for more background on the Frobenius method.
@@
In [FLPW24, Conjecture 1.2], the authors formulate a precise conjecture for the limiting connection probabilities of general critical loop $O(n)$ models, which also give the conjectural connection probabilities for multichordal $\operatorname{CLE}_{\kappa}$ with $2N$ boundary arcs (see [AMY25]). In the regime $\kappa\in(4,8)$, this conjecture was proved for $n=\sqrt{2}$ (FK-Ising) [FPW24] and for $2N=4$ [MW18]. In principle, by taking $2N=8$ and shrinking four of the eight boundary arcs, one could obtain the CLE boundary Green’s functions (1.3) via renormalized limits of the corresponding conjectural connection probabilities. However, this does not seem practically feasible, since the explicit formulas for these conjectural connection probabilities involve the pure partition functions, whose constructions are highly nontrivial [FK15, KP20, KP16, Wu20, Zha24, FLPW24]. It is therefore unclear whether our results can be recovered in this way from these conjectural formulas.
@@ 1.3 Discussions
We now briefly discuss our proof strategy. The first step is to identify a precise correspondence between $\operatorname{CLE}_{\kappa}$ and the $\operatorname{SLE}_{\kappa}$ bubble measure for $\kappa\in(4,8)$. To achieve this, we show that the loop sampled from the counting measure on boundary-touching loops in a $\operatorname{CLE}_{\kappa}$ configuration is equal in law to the unrooted $\operatorname{SLE}_{\kappa}$ bubble measure, see Proposition $2.$ This identification allows us to express the boundary Green’s functions of $\operatorname{CLE}_{\kappa}$ (1.3) in terms of those of the $\operatorname{SLE}_{\kappa}$ bubble measure. Furthermore, $\operatorname{SLE}_{\kappa}$ bubble measure is also the weak limit of the chordal $\operatorname{SLE}_{\kappa}$ when its starting point approaches its end point. According to [FZ23], the boundary Green’s functions of chordal $\operatorname{SLE}_{\kappa}$ are finite and satisfy a martingale property as the chordal $\operatorname{SLE}_{\kappa}$ grows. Combined with their smoothness, the martingale property then yields a pair of second-order PDEs, known as BPZ equations. Establishing such smoothness is nontrivial and relies on Hörmander’s hypoellipticity; see Lemma 2.8, as inspired by [Dub15b, FZ23]. Then, we apply a fusion procedure to these PDEs to derive a third-order differential equation satisfied by boundary Green’s functions of the $\operatorname{SLE}_{\kappa}$ bubble, which corresponds to collapsing two marked boundary points. As mentioned before, this step follows from the framework in [Dub15a], which also appeared in recent works [KLPR25, LPR25] for specific values of $\kappa$. This proves Theorem 1.3.
@@
Acknowledgment. The author thanks Xin Sun for his suggestions on the draft of this paper. This work was supported by the National Natural Science Foundation of China (Grant No. 12526204) and National Key R&D Program of China (No. 2021YFA1002700).
@@ 2 SLE bubble measures and CLE boundary Green’s functions
We first review the background on the SLE bubble measure in Section 2.1, and then show the relation between boundary-touching CLE loops and the (unrooted) SLE bubble measure in Section $2.$ In Section 2.3, we relate the CLE boundary four-point connectivities (1.3) to the boundary Green’s functions of SLE bubble and of chordal SLE. In Section 2.4, we show the smoothness of these boundary Green’s functions, thus obtaining a pair of second-order PDEs satisfied by the boundary Green’s functions of chordal SLE, which will be the basis of the fusion procedure in Section 3.
@@ 2.2 Boundary-touching CLE loops and unrooted SLE bubble measure
Now, since $\rho$ is obtained from the counting measure on boundary-touching loops and every $z\in\mathbbm{H}$ is surrounded by at most one CLE loop, we have ${\bf 1}_{\ell\text{ surrounds }z}\rho(d\ell)dz={\bf 1}_{\ell\cap\mathbbm{R}\neq\emptyset}\Theta_{z}(d\ell)dz$, where dzdz is the Lebesgue measure on $\mathbbm{H}$. Combined with the above result, we find ${\bf 1}_{\gamma\text{ surrounds }z}\mu^{\mathrm{Bub}}(d\gamma)dz=C{\bf 1}_{\ell\text{ surrounds }z}\rho(d\ell)dz$. The result then follows by deweighting the Lebesgue areas of the regions surrounded by the bubble (resp. loop) on both sides. ∎
@@ 2.3 CLE boundary Green’s functions
The aim of this section is to relate the $\operatorname{CLE}_{\kappa}$ boundary Green’s functions defined in (1.3) to the limit of boundary Green’s function of chordal $\operatorname{SLE}_{\kappa}$; see Proposition 2.7. Due to symmetry, it suffices to focus on $G^{(1234)}(x_{1},x_{2},x_{3},x_{4})$ and $G^{(12)(34)}(x_{1},x_{2},x_{3},x_{4})$. We refer readers to Appendix A for further background of $\operatorname{SLE}_{\kappa}$ boundary Green’s functions, including the various definitions appeared in the literature [Law15, Zha22, FZ23] and their equivalence to the definition in this paper.
@@
Let $\mathfrak{C}$} be the same constant as in Proposition $2.$ Then for any $n\geq 2$, we have
@@
where the integration is taken over $\mu_{\mathbbm{H},y_{1},y_{2}}^{\#}(d\eta)$. (2.12) is equivalent to the boundary Green’s function considered in [Law15, FZ23]; see Proposition A.2.
@@ 2.4 Smoothness and second-order PDEs
Note that the Lie brackets ${\bf X},{\bf Y}]=-\sum_{i=1}^{3}\frac{1}{x_{i}^{2}}\partial_{x_{i}}$ and ${\bf X},[{\bf X},{\bf Y}]]=2\sum_{i=1}^{3}\frac{1}{x_{i}^{3}}\partial_{x_{i}}$. Since $x_{i}$’s are mutually different on $U$, the Lie algebra generated by $\bf X$} and $\bf Y$} has rank 33 on $U$. Thus, $\mathcal{L}$ satisfies the Hörmander’s condition222i.e. for two vector fields $\bf X$} and $\bf Y$}, ${\bf X},{\bf Y}$ and their iterated Lie brackets ${\bf X},{\bf Y}],[{\bf X},[{\bf X},{\bf Y}]]$ etc. span the whole tangent space at every point.. By [Bon69, Theorem 5.2], we can then define a Poisson operator $\mathcal{P}:C(\partial U)\to C^{\infty}(U)\cap C(\overline{U})$ for $\mathcal{L}$ such that for any $\omega\in C(\partial U)$, $(\mathcal{P}\omega)|_{\partial U}=\omega$ and $\mathcal{L}(\mathcal{P}\omega)=0$.
@@ 4 Identification of solutions
Here to the right side we use Lemma $4.$ Combined with $r_{3}^{-h}r_{4}^{-h}|\mathsf{m}_{r_{3},r_{4}}|\to G_{\mathbbm{H},x_{1},x_{2}}(x_{3},x_{4})$, the result then follows. ∎
@@
where $\tau$ and $\phi$ are defined in Corollary 4.6, and
@@
Here and after, $C>0$ is some constant depending on $x_{1},x_{3},x_{4}$ and can vary from line to line. To see (4.14), first note that for any compact hull AA, $A\cap\mathbbm{R}=[a,b]$, $x>b$ such that ${\rm dist}(x,A)>d_{1}>0$, for the conformal map $\psi:\mathbbm{H}\setminus A\to\mathbbm{H}$ with $\psi(a)=0$, $\psi(b)=\infty$ and $\psi(x)=1$, we have $|\psi^{\prime}(x)|\leq 4d_{1}^{-1}$ by Koebe’s 1/4 theorem. By conformal covariance, this implies the boundary one-point Green’s function $G_{\mathbbm{H}\setminus A,a,b}(x)\leq Cd_{1}^{-h}$. Now suppose $b<y<x$, ${\rm dist}(y,A)>d_{2}>0$ and $|x-y|>d_{1}+d_{2}$. According to the martingale property of the boundary two-point Green’s function (see [FZ23, Theorem 4.1]), we have
@@
here $\mathbbm{E}_{x}$ is with respect to the conditional probability measure $\mathbbm{P}_{x}$ of the chordal $\operatorname{SLE}_{\kappa}$ curve $\eta$ on $\mathbbm{H}\setminus A$ from $a$ to $b$ conditioned to hit $x$ (we denote this hitting time by $\sigma$), and $U^{*}$ is the connected component of $\mathbbm{H}\setminus(A\cup\eta[0,\sigma])$ such that $b$ is on its boundary. Furthermore, following the estimate in [FZ23, Lemma 3.5], for $s\in(0,1)$, the probability that a chordal $\operatorname{SLE}_{\kappa}$ on $\mathbbm{H}$ from 0 to $\infty$ hits $\partial B(1,\frac{1}{10})$ after hitting $B(1,\frac{1}{10}s)$ and before disconnecting 11 from $\infty$ is bounded above by $Cs^{2h}$. The domain Markov property of $\operatorname{SLE}_{\kappa}$ (see e.g. [FZ23, Eq.(3.27)]) then implies $\mathbbm{P}_{x}[{\rm dist}(U^{*},y)<rd]\leq Cr^{2h}$ for $r\in(0,\frac{1}{10})$. Consequently, we obtain $G_{\mathbbm{H}\setminus A,a,b}(x,y)\leq Cd_{1}^{-h}d_{2}^{-h}$, thus proving (4.14).
@@
Then (4.17) has three linearly independent solutions $V_{0}(\lambda)$, $V_{2}(\lambda)$ and $V_{\frac{1}{3}}(\lambda)$, which are defined in Section $1.$ See also [GV18, Appendix C]. In particular, $V_{0}(\lambda)=V_{0}(1-\lambda)$, and as $\lambda\to 0$,
@@
    They correspond to $G^{(14)(23)}$ and $G^{(12)(34)}$, respectively, and have the following probabilistic interpretation. Let $e_{1}$ (resp. $e_{2}$) be the union of two independent Brownian excursions from 0 to $\infty$ (resp. from $\lambda$ to 11) such that $e_{1},e_{2}$ are also independent. Let $P(\lambda)$ be the probability that $e_{1}$ does not intersect with $e_{2}$. Then we have $P(\lambda)\propto\frac{U_{1}(\lambda)}{\lambda^{4}}$, i.e.
@@
    This expression can be proved similarly to Section 3, using the fact that the right boundary of $e_{1}$, which has the law of $\operatorname{SLE}_{8/3}(2)$ (see [LSW03]), is the fusion limit of a pair of $\operatorname{SLE}_{8/3}$ from 0 to $\infty$, as well as the half-plane intersection exponent $P(\lambda)=\lambda^{3+o(1)}$ as $\lambda\to 0$ [LSW01].
@@ 5 One-Bulk and two-boundary connectivities
By symmetry, the above PDE also holds when changing the place of $y_{1}$ and $y_{2}$. Here for convenience, we view $z$ and $\bar{z}$ as independent variables. Note that such equations are similar to those in Proposition 2.9, with the three marked points $(x_{1},h),(x_{2},h),(x_{3},h)$ replaced by $(x,h),(z,\frac{\alpha}{2}),(\bar{z},\frac{\alpha}{2})$. Consequently, we can repeat the fusion procedure in Section 3 in parallel. Namely, (5.2) implies a third-order differential equation for $G(u,x,z)$. The resulting ODE [^5] Since $\lambda\in\mathbbm{C}$ takes value on the circle $\{\lambda\in\mathbbm{C}:|\lambda-1|=1\}$, we can view $\lambda$ as if it were a real variable. for $\Delta(\lambda)$ is